# Make Some Noise: Unlocking Language Model Parallel Inference Capability through Noisy Training

### v1

@@ Abstract
Yixuan Wang1, [^2] Equal contribution., Xianzhen Luo1, [^2] Equal contribution., Fuxuan Wei1, Yijun Liu1, Qingfu Zhu1, Xuanyu Zhang2, Qing Yang2, Dongliang Xu2, Wanxiang Che1, [^1] Corresponding author., 1 Harbin Institute of Technology, Harbin, China 2 Du Xiaoman (Beijing) Science Technology Co., Ltd. {wyx,xzluo,fxwei,yjliu,qfzhu,car}@ir.hit.edu.cn {zhangxuanyu,yangqing,xudongliang}@duxiaoman.com
@@ 2.1 Jacobi Decoding
$\begin{cases}y_{1}&=\arg\max P_{\theta}(y_{1}|x)\\ y_{2}&=\arg\max P_{\theta}(y_{1}|y_{1},x)\\ &\vdots\\ y_{m}&=\arg\max P_{\theta}(y_{m}|y_{1:m-1},x)\\ \end{cases}$ (1)
@@ 2.2 Speculative Decoding
In addition, there are speculative decoding methods that do not require training. Fu et al. (2024) performs more efficient verification by collecting n-gram segments generated during Jacobi decoding as draft tokens. Saxena (2023) achieves acceleration in specific domains simply by retrieving draft tokens from the ahead prompt. He et al. (2023) enable plug-in draft token generation by retrieving a constructed knowledge database.
@@ 4.1 Experimental Setup
Methods Metric(+) Speed Speedup
@@
Table 1: Results of task performance experiments in general and code domains. The general domain metric uses scores from MT-bench and the code domain uses pass@1 under greedy decoding. For the code domain, we choose the average of HumanEval and MBPP as a composite metric. ‘(+)’: Results after executing additional tests from evalplus. $\uparrow$’: Percentage improvement over models without MSN.
Methods AS MT-B Trans Sum QA Math RAG #MAT #Speed Overall
@@
In this paper, we conduct experiments on the task performance and acceleration performance of MSN, respectively. For task performance, we use the MT-bench Zheng et al. (2024) in the general domain, the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks in the code domain for evaluation. Evalplus Liu et al. (2023) , which provides additional test cases for problems in HumanEval and MBPP, is also included. For acceleration performance, we performed a speedup evaluation of the proposed parallel decoding methods on Spec-Bench Xia et al. (2024). This benchmark contains data from multiple domains and provides a fair comparison with existing acceleration methods. Following previous work, all speed related experiments are done on a single A100-80G device with the batch size as 1. For our MSN model, the draft token length during inference is consistent with the noise segment length during training, which is 4.
@@ Results.
The metric in Table 1 represents the task performance of each model. There is no significant performance loss of the model trained by MSN on the downstream task compared to SFT. Futhermore, The MSN model even delivers a slight performance boost in both domain. The enhancement in the code domain is particularly noteworthy, given that evaluating generated programs is more rigorous than evaluating conversation. Programs must be correctly formatted and pass all test cases to be deemed successful. It indicates that MSN does not hurt the model to acquire capabilities during the SFT phase. Our analysis suggests that this gain comes from the fact that noise mitigates the negative effects of teacher forcing training on the model to some extent. The causal denoising task forces the model to focus on more distant tokens when predicting the current location token because the current input is noisy.

### v2

@@ Abstract
Yixuan Wang1, [^2] Equal contribution., Xianzhen Luo1, [^2] Equal contribution., Fuxuan Wei1, Yijun Liu1, Qingfu Zhu1, [^1] Corresponding author., Xuanyu Zhang2, Qing Yang2, Dongliang Xu2, Wanxiang Che1 1 Harbin Institute of Technology, Harbin, China 2 Du Xiaoman (Beijing) Science Technology Co., Ltd. {wyx,xzluo,fxwei,yjliu,qfzhu,car}@ir.hit.edu.cn {zhangxuanyu,yangqing,xudongliang}@duxiaoman.com
@@ 2.1 Jacobi Decoding
$\begin{cases}y_{1}&=\arg\max P_{\theta}(y_{1}|x)\\ y_{2}&=\arg\max P_{\theta}(y_{2}|y_{1},x)\\ &\vdots\\ y_{m}&=\arg\max P_{\theta}(y_{m}|y_{1:m-1},x)\\ \end{cases}$ (1)
@@ 2.2 Speculative Decoding
In addition, there are some speculative decoding methods that do not require training. LLMA Yang et al. (2023) achieves 2$\sim$3x speedups on tasks such as conversations by retrieving text segments from reference texts. Fu et al. (2024) performs more efficient verification by collecting n-gram segments generated during Jacobi decoding as draft tokens. Saxena (2023) achieves acceleration in specific domains simply by retrieving draft tokens from the ahead prompt. REST He et al. (2023) enables plug-in draft token generation by retrieving a constructed knowledge database. Zhao et al. (2024) proposes Ouroboros that combines the advantages of both the retrieval and draft model approaches. It utilizes the retrieval method to further enhance the generation length of the draft model, achieving a significant speedup ratio.
@@ 4.1 Experimental Setup
Method Metric(+) Speed Speedup
@@
Table 1: Results of task performance experiments in general and code domains. The general domain metric uses scores from MT-bench and MMLU-weighted from MMLU. The code domain uses pass@1 under greedy decoding. For the code domain, we choose the average of HumanEval and MBPP as a composite metric. ‘(+)’: Results after executing additional tests from evalplus. $\uparrow$’: Percentage improvement over models without MSN.
Method AS MT-B Trans Sum QA Math RAG #MAT #Speed Overall
@@
In this paper, we conduct experiments on the task performance and acceleration performance of MSN, respectively. For task performance, we use the MT-bench Zheng et al. (2024) and MMLU Hendrycks et al. (2020) in the general domain, the HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks in the code domain for evaluation. Evalplus Liu et al. (2023) , which provides additional test cases for problems in HumanEval and MBPP, is also included. For acceleration performance, we performed a speedup evaluation of the proposed parallel decoding methods on Spec-Bench Xia et al. (2024). This benchmark contains data from multiple domains and provides a fair comparison with existing acceleration methods. Following previous work, all speed related experiments are done on a single A100-80G device with the batch size as 1. For our MSN model, the draft token length during inference is consistent with the noise segment length during training, which is 4.
@@ Results.
The metric in Table 1 represents the task performance of each model. There is no significant performance loss of the model trained by MSN on the downstream task compared to SFT. Futhermore, The MSN model even delivers a slight performance boost in both domain. The enhancement in the code domain is particularly noteworthy, given that evaluating generated programs is more rigorous than evaluating conversation. Programs must be correctly formatted and pass all test cases to be deemed successful. It indicates that MSN does not hurt the model to acquire capabilities during the SFT phase. Our analysis suggests that this gain comes from the fact that noise mitigates the negative effects of teacher forcing training on the model to some extent. The causal denoising task forces the model to focus on more distant tokens when predicting the current location token because the current input is noisy. We also conduct som experiments comparing SFT and MSN with different base models in the code domain, which can be found in Appendix C.
@@ 6 Conclusion
Acknowledgements
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC) via grant 62236004, 62206078, 62441603 and 62476073 and the support of Du Xiaoman (Beijing) Science Technology Co., Ltd.
@@ Appendix B Templates for Token Tree
Appendix C More Experiment Results
In addition to the previously mentioned DeepSeek-Coder, we have experimented on a variety of different code base models (CodeLlama Roziere et al. (2023), StarCoder Li et al. (2023)). The results of the experiments on Humaneval and MBPP are shown in Table 5. We can see that the MSN and SFT methods do not reflect a significant gap on models of different sizes and sources.
BaseModel Method Humaneval Humaneval+ MBPP MBPP+
CodeLlama 7B SFT 62.20 57.90 65.60 57.70
CodeLlama 7B MSN 64.00 57.90 63.50 54.50
StarCoder 3B SFT 57.30 53.70 60.60 52.40
StarCoder 3B MSN 59.10 55.50 60.30 52.60
StarCoder 7B SFT 68.90 64.60 63.80 55.00
StarCoder 7B MSN 66.50 60.40 64.30 55.60
Table 5: Experimental results for different code base models.
Method humanities stem social-science other MMLU MMLU-weighted
SFT 65.18 51.74 69.72 64.71 61.55 60.38
MSN 66.32 51.20 71.09 64.33 61.83 60.87
Table 6: Complete experiment results on MMLU Benchmark. Both SFT and MSN methods are trained on Llama3-8B-Base model.